# Benchmarking LLM-based Machine Translation on Cultural Awareness

### v2

@@ Abstract
Translating cultural-specific content is crucial for effective cross-cultural communication. However, many MT systems still struggle to translate sentences containing cultural-specific entities accurately and understandably. Recent advancements in in-context learning utilize lightweight prompts to guide large language models (LLMs) in machine translation tasks. Nevertheless, the effectiveness of this approach in enhancing machine translation with cultural awareness remains uncertain. To address this gap, we introduce a new data curation pipeline to construct a culturally relevant parallel corpus, enriched with annotations of cultural-specific items [^1] The code and data will be released at: https://github.com/BigBinnie/Benchmarking-LLM-based-Machine-Translation-on-Cultural-Awareness. Furthermore, we devise a novel evaluation metric to assess the understandability of translations in a reference-free manner by GPT-4. We evaluate a variety of neural machine translation (NMT) and LLM-based MT systems using our dataset. Additionally, we propose several prompting strategies for LLMs to incorporate external and internal cultural knowledge into the translation process. Our results demonstrate that eliciting explanations can significantly enhance the understandability of cultural-specific entities, especially those without well-known translations.
@@ 1 Introduction
Despite the notable success of MT systems in recent years, thanks to multilingual pre-trained language models Aharoni et al. (2019), their performance when translating cultural-specific items (CSIs) remains poor. This is primarily due to the gap between the cultural differences associated with languages Akinade et al. (2023); Liebling et al. (2022). The variation of cultural-specific items across cultures Woolford (1983) has been a long-standing aspect of translation studies Newmark (1988); Persson (2015); Fernández Guerra (2012). However, existing terminology translation datasets and methods mostly focus on popular domains (e.g., medicine, finance) Dinu et al. (2019); Ghazvininejad et al. (2023), yet do not cover much on cultural-specific aspects. Distinguished from general terms, cultural-specific items are unique to particular cultural groups, making their literal translations difficult for people from other cultures to understand. Moreover, many CSIs have no translations into other languages, further increasing the difficulty of data collection and evaluation of the translation performance. For example, in Figure 1, a Taiwanese dish called “五更肠旺”, which is chitterling in hot pot, has no well-known existing English translations. The translation of ChatGPT “Wu Geng Intestine Soup” is misleading and can not be easily understood by English native speakers. In contrast, a professional human translator would elicit an explanation to make the translation more understandable for readers. As a result, the nuanced characteristics of CSIs and the limited availability of data sources pose challenges in building culture-aware MT systems that can translate these terms into understandable ones in target languages.
Figure 1: Cultural translation errors made by Google Translate and ChatGPT systems
Recently, a new translation paradigm has emerged, which employs prompts to guide large language models (LLMs) to perform machine translation Brown et al. (2020) in a zero-shot or few-shot fashion. With this flexible paradigm, cultural knowledge can be seamlessly incorporated into LLM translation prompts. However, LLM-based translation is sensitive to prompting strategies and prone to generating hallucinations Ji et al. (2023). As shown in Figure 1, the translation of “A Few Good Men” as “壮志凌云” (i.e., Top Gun) wrongly refers to a completely different film. Given that, the cultural sensitivity of LLM-based MT systems remains an open question, especially when dealing with culturally relevant content. Moreover, there is scant research comparing LLM-based translation versus NMT systems regarding their cultural awareness, partly due to the absence of a rich-annotated culturally sensitive parallel corpus and reliable automated metrics for evaluation. This scarcity further hinders the MT development in navigating cultural nuances, limiting the MT systems’ efficacy in promoting cross-cultural communication.
In this study, we focus on evaluating the culture-awareness of MT systems, addressing two research questions: (1) NMT v.s. LLM-based MT, which group of MT systems embed more cultural awareness? (2) Diving into LLM-based MT, what prompting strategy benefits the most for guiding LLMs in cultural-specific translation? To this end, we propose a novel data curation pipeline to build a culture-centered parallel corpus with limited human annotation efforts. To capture challenging examples with geo-metadata, we further conduct cultural knowledge augmentation for CSIs. Our parallel corpora include 6 language pairs (i.e., English-Chinese, English-French, English-Spanish, English-Hindi, English-Tamil, and English-Telugu), covering 7,253 CSIs in 18 concept categories from over 140 countries and regions.
To assess the cultural nuance, we design new automatic metrics targeting the translation quality of cultural concepts, both on the translation accuracy and understandability. Different from existing evaluation metrics which mainly focus on semantic similarity with the reference Anastasopoulos et al. (2021), we define a new metric called understandability to evaluate the CSI translation in a reference-free manner. We use our metrics to compare a series of MT systems, including LLM-based MT, traditional NMTs, and commercial MT systems, on translating cultural content. Moreover, we examine a set of prompting strategies to endow cultural knowledge into LLM-based MT. The traditional terminology translation method enhances the accuracy of translating CSIs that have well-known references but falls short of translating no-translation CSIs. However, incorporating CSI explanations in the prompt significantly improves the translation quality, especially for no-translation CSIs. Our further investigation into different MT systems’ translation strategies reveals that employing more sophisticated approaches, such as describing CSIs and substituting them with similar content in the target language, effectively improves the readability of CSI translations.
2 Culturally Relevant Data Construction
Figure 2: A data curation pipeline for constructing a cultural-specific English-Chinese parallel corpus.
A wide range of well-structured multilingual knowledge gathered by various crowd workers in Wikipedia makes it an ideal source for collecting CSIs and their descriptions in aligned languages. The overall workflow of our data collection includes four stages: (1) building a wiki-centered cultural taxonomy (§2.1); (2) curating parallel entities related to diverse sociocultures (§2.2); (3) augmenting geo-metadata (§2.3). The curation pipeline with one example is shown in Figure 2. We have more data examples in Appendix A.
@@ 2.1 Cultural Taxonomy Extraction
Given that culture is an abstract concept, it is hard to directly capture fine-level cultural characteristics from texts. With this consideration in mind, we referred to an existing CSI classification framework  Newmark (1988), which has been popularly used in the study of human translations of cultural concepts, to identify culturally relevant texts from Wikipedia. Specifically, there are five categories in this framework, including: 1) ecology; 2) material culture; 3) social culture; 4) organizations, customs, ideas; 5) gestures and habits. Following a primary investigation of Wikipedia texts, we find that each entity-centered Wikipedia page is labeled by a variety of Wikipedia categories. To save the efforts of matching each entity on Wikipedia with each CSI category, we turn to map CSI categories with Wikipedia categories  Asthana and Halfaker (2018). To further guarantee the mapping quality, we manually create a mapping table for building connections on the categories from two resources. The resulting mapping table and category classification tool are described in Appendix 3.
2.2 Cultural Parallel Text Collection
To construct a cultural parallel text corpus (e.g., for En-Zh), we collect public text articles from Wikipedia’s translation tool that cover a wide range of cultural topics, and conduct sentence alignment to get a parallel corpus from that (The data collection details are in Appendix §C). To expand the language coverage in our corpora, we also reuse open-source parallel datasets from OPUS Tiedemann (2016). These include Wikipediav1.0 for English-French and English-Spanish, as well as Samanantarv0.2 for English-Hindi, English-Tamil, and English-Telugu. To identify cultural-specific sentences, we perform entity-linking  Ringgaard et al. (2017) to identify Wikipedia entities on the source texts, and use Wikiproject classification tool Asthana and Halfaker (2018) to classify these entities into cultural categories which are further mapped to our CSI categories using the cultural taxonomy (§2.1). Finally, we only keep the texts that contain items belonging to CSI categories.
@@ 2.3 Cultural Knowledge Augmentation
Existing MT studies Arthur et al. (2016); Hu et al. (2022) have been using external knowledge sources (e.g., Wikidata) to improve named entity translations. To enable future adaptations of these studies on our collected corpus, we parse Wikidata to extract the metadata of CSIs, which include their cultural labels, descriptions, and aliases in multiple languages. Furthermore, to identify items which are culture-specific, we collect information on the country of origin for each items and remove sentences containing items that do not have an associated origin country. This meticulous approach enabled us to enrich our dataset with supplementary information that can be utilized to evaluate the performance of machine translation models when handling culturally specific content.
2.4 Data Characteristic Summary
@@
Table 1 shows the statistics of our parallel corpora for the evaluation of MT systems on six language pairs. Particularly, for each language pair, we count the total number of detected CSIs by CSIs Counts and the number of unique CSIs by CSIs Types. It’s noted that not all the CSIs have translations on Wikidata, so we the number of CSIs containing the translations in WikiData by CSI Translations. Considering that many CSIs only exist within a specific culture group, which can’t be located in the parallel corpus, CSIs which don’t have a translation in other languages should take a higher proportion in real-world corpus than in our dataset.
@@ 3 Cultural Awareness Evaluation
Existing evaluation methods for terminology machine translation primarily focused on assessing the adequacy and fluency of the translated terms Anastasopoulos et al. (2021). However, CSI usually has more than one translation using various strategies, leading to challenges of fine-grained evaluation.
@@
To evaluate the accuracy of translations, we introduce the CSI-Match metric as a modification to the traditional Exact Match (EM) evaluation metric. CSI-Match measures the accuracy of term translation using a more nuanced, fuzzy matching approach. CSI-Match is the maximal partial similarity ratio (PSR) between the reference CSI translations $\{t_{1},t_{2},\ldots,t_{n}\}$ and the system output sentence $S$. CSI-Match is calculated by Eq. (1), resulting in a value from 00 to $100$. A higher value indicates a stronger similarity of the predicted CSIs over a set of translation references.
@@
where $S_{i:j}$ is a substring of $S$ from word index $i$ to $j$, $d(,)$ is the normalized Levenshtein distance Levenshtein et al. (1966) between two strings.
Understandability in Human Evaluation:
In the context of evaluating CSI translation, we believe that, in addition to adequacy and fluency, the translated terms should also be comprehensible to speakers of the target language, considering that certain terms may not exist in the target language group. So, we design a human evaluation targeting at understandability which we define as the degree to which native target-language speakers can understand and explain the CSI translations. Specifically in the human evaluation, we ask the native target-language speaker to rank the understandability of CSI generated by various MT systems and evaluate the quality of CSI translation accordingly. To enhance our metric’s practical application, we leverage GPT-4 to compare the understandability between the target and MT systems, which has been shown as an effective way to evaluate the generation performances Rafailov et al. (2023). We calculate the understandability of each MT system as the win rate of the comparison. The evaluation prompt is detailed in Appendix §F.
Translation Strategies:
Furthermore, we categorize the translation strategies of CSI based on prior translation theories Newmark (1988); Persson (2015). These theories define different categorizations of strategies to improve the understandability of CSIs while keeping the cultural specialty. We select 4 strategies that are common to our dataset which are explained in Table 2. We then conduct a human evaluation to annotate the translation strategies employed by different MT models, and compare them with the strategies used for generating the reference translations.
Strategy Explanation
Transliteration Phonetically translates source CSIs
Literal translation Directly translates word-by-word
Description Integrates CSI descriptions of the CSIs
into the translation
Substitution Replace source CSIs by a semantically
equivalent item in the target language.
Table 2: Translation Strategies
4 Cultural Knowledge Prompting
In this section, we explore various prompting strategies to introduce cultural knowledge for LLM MT. We elucidate our strategies for generating in-context examples from external knowledge, which involve employing CSI translation pairs and CSI explanations. Additionally, we delve into several prompting strategies eliciting from LLMs’ internal knowledge. Table 3 shows examples of various prompting strategies.
@@ 4 Cultural Knowledge Prompting
cannoli:里考塔芝士卷(Ricotta cheese rolls), 奶油甜馅煎饼卷 (Sweet Cream pancake rolls)
@@
Self-Ranking (SR) User: Translate the following English text to Chinese: [Source Sentence].
Please give [Generated Number] most likely translations, and ensure ”cannoli” in each result to correspond to different translations
LLM: [Translations]
User: Please select the best translation result of entities: cannoli. The translation of entitis in the sentence
should be the closest to its explanation, and is easy to be understood by Chinese readers.
@@
Knowledge Translations:cannoli:里考塔芝士卷(Ricotta cheese rolls), 奶油甜馅煎饼卷 (Sweet Cream pancake rolls)
@@
Table 3: Prompting strategy examples (Top) and a source with cultural knowledge for En-Zh translation (Bottom)
4.1 External CSI Translation (CT)
We examine an external CSI dictionary for improving LLM-based MT. In particular, bilingual translation dictionaries play a vital role in the workflow of human translators Ghazvininejad et al. (2023). Here, we assess the impact of incorporating a CSI dictionary within the prompts. Specifically, we incorporate CSIs along with their corresponding translations prior to a basic translation instruction.
4.2 External CSI Explanation (CE)
CSIs may not have a direct equivalent in the target language’s culture when the concepts are not commonly used by the target language speakers. Therefore, it becomes necessary to translate based on the explanation of the CSI to assist the target audience in better understanding the content. To assess the impact of explanations, we include the CSI description obtained from Wikipedia in the prompt before the basic translation instructions. This allows us to study whether additional explanations of CSIs can enhance the MT performance.
4.3 Self-Explanation (SE)
We also examine LLMs’ internal knowledge for explaining the meaning of CSIs. Notably, chain-of-Thought (CoT) prompting has shown effective in eliciting LLMs’ internal knowledge Wei et al. (2022); Kojima et al. (2022). Inspired by this, we treat the explanation of CSIs in a source sentence as intermediate reasoning steps before translating the whole sentence. We then design the explanation prompting strategy in two steps for machine translation. In the first step, we prompt the LLM to explain the meaning of all CSIs in the source sentence. In the second step, we ask the LLM to translate the whole sentence by combining the LLM’s explanation with another prompt instruction.
4.4 Self-Ranking (SR)
Finally, LLMs can be prompted to sample different translation outputs. Even for multiple semantically equivalent prompts, the word choice in the prompts can have a significant impact on the performance of LLM models in translation tasks. To achieve more consistent and reliable results, we examine a self-ranking method Wu et al. (2023) to instruct LLMs to rank their translation outputs. Specifically, we prompt the model to first generate a fixed number of potential translations.

### v3

# Benchmarking Machine Translation with Cultural Awareness

@@ Abstract
Translating culture-related content is vital for effective cross-cultural communication. However, many culture-specific items (CSIs) often lack viable translations across languages, making it challenging to collect high-quality, diverse parallel corpora with CSI annotations. This difficulty hinders the analysis of cultural awareness of machine translation (MT) systems, including traditional neural MT and the emerging MT paradigm using large language models (LLM). To address this gap, we introduce a novel parallel corpus, enriched with CSI annotations in 6 language pairs for investigating Culturally-Aware Machine Translation—CAMT. [^1] The corpus and code are released at https://github.com/BigBinnie/Benchmarking-LLM-based-Machine-Translation-on-Cultural-Awareness Furthermore, we design two evaluation metrics to assess CSI translations, focusing on their pragmatic translation quality. Our findings show the superior ability of LLMs over neural MTs in leveraging external cultural knowledge for translating CSIs, especially those lacking translations in the target culture.
@@ 1 Introduction
Machine translation (MT) systems have achieved remarkable success in recent years, thanks in part to the pre-trained backbones of multilingual language models Aharoni et al. (2019) and the availability of multilingual corpora NLLB Team et al. (2022). Despite these advances, terminology translation remains challenging in both general contexts Dinu et al. (2019) and specific domains like medicine and law Ghazvininejad et al. (2023). Many existing MT studies on terminology translation have focused on breaking language barriers rather than cultural barriers, often assuming that literal (i.e., word-for-word) translation pairs already exist for the common knowledge shared by speakers of both the source and target languages Anastasopoulos et al. (2021). However, culture is deeply intrinsic to language, and language translation entails cross-cultural communication  Newmark (1988); Fernández Guerra (2012). Due to the diverse nature of knowledge and norms across cultures, many cultural-specific items (CSIs) related to food, clothing, art, and religion are rarely used by speakers outside the items’ associated cultural group, with some items even not existing in certain cultures Woolford (1983); Persson (2015). As a result, cultural-specific items usually lack available literal translations across languages, leading most MT systems to perform poorly on cultural-centered translations in real-world deployment  Akinade et al. (2023); Liebling et al. (2022). As shown in Figure 1, common errors, including copy and factual errors, are still made by the state-of-the-art MT systems (e.g., Google Translate and ChatGPT). More importantly, the nature of CSIs leads to difficulty in collecting high-quality, yet diverse parallel corpora at scale, hindering a systematic evaluation of both the traditional neural MT systems and the emerging MT paradigm using large language models (LLM).
Figure 1: Culture-specific item translation errors.
To address this challenge, a handful of recent studies have begun to curate culture-related corpora for analysis from two main perspectives. The first perspective emphasizes regional varieties, such as the variety of Portuguese used in Brazil versus Portugal Riley et al. (2023). The second perspective focuses on cultural content in translation, such as recipes Cao et al. (2024) and idioms Li et al. (2024). Given the difficulty of obtaining word-for-word translation pairs for CSIs, these studies are shifting from the traditional literal translation paradigm to free translation which aims to convey the meaning of source texts and prioritizes readability and cultural relevance over strict accuracy and structural fidelity. Despite the valuable contributions of these works, two major concerns may still hinder the analysis of MT systems in navigating cultural nuances. First, the demand for high-quality data requires costly human annotations, restricting these studies from scaling up their curated data resources in terms of size, language pair diversity, and cultural domain coverage (see Table 1). Second, common MT evaluation metrics designed for literal translation lead to a lack of reliable assessments of free translation quality.
In this study, we address the two aforementioned concerns with a particular focus on translating culture-specific items. Specifically, we introduce an annotation-efficient data curation pipeline that can freely gather a diverse, large-scale, culture-centered parallel corpus while ensuring data quality. The resulting corpus, called Culturally-Aware Machine Translation (CAMT), encompasses 6 language pairs, covering 6,983 CSIs across 18 concept categories from 235 countries and regions. To facilitate automatic assessment emphasizing CSI-centered free translation, we propose two evaluation metrics: CSI-Match and pragmatic translation assessment (PTA). The CSI-Match metric evaluates the translation accuracy of isolated CSIs, independent of their sentence context. In contrast, the PTA metric assesses the comprehensibility of CSI translations within sentence contexts, emphasizing their pragmatic effectiveness in communication with native speakers of the target language.
Leveraging our CAMT corpus and designed evaluation metrics, we conduct a systematic analysis to investigate the capability of state-of-the-art neural MT and LLM-based MT systems in translating cultural content. First, we examine their efficacy with two popular terminology translation strategies that utilize the CSI dictionary. Our findings indicate that the terminology translation strategies greatly enhance the CSI translation accuracy for both neural MT and LLM-based MT systems. However, LLMs exhibit superior capability in leveraging the external dictionary compared to NMTs, particularly for CSIs that lack well-known translations. Next, to further examine LLMs’ capability for integrating external knowledge into translations, we explore prompting strategies that incorporate the CSI explanations in the prompts. Our results show that incorporating CSI explanations in the prompts notably improves the pragmatic translation quality, especially for CSIs without direct translations. In summary, our contributions are as follows:
  • We curate a diverse parallel corpus in six language pairs with rich cultural-specific item annotations using a highly automatic pipeline.
  • We introduce two new evaluation metrics (CSI-Match and PTA) to assess translation quality regarding cultural nuances, particularly for terms lacking established translations.
  • We examine both LLM-based MT and NMT systems using our dataset and metrics. Our results indicate that LLMs can effectively incorporate external cultural knowledge, thereby improving the pragmatic translation quality of CSIs.
2 Related Works
Culturally-Aware Machine Translation:
As languages and cultures are highly intertwined, there is a growing desire to empower cultural awareness of MT systems Hershcovich et al. (2022); Riley et al. (2023). However, as cultural nuances are subtle, collecting culturally sensitive data Akinade et al. (2023) remains costly and time-consuming. Therefore, current work on cultural-aware translations is limited to specific domains and language pairs Cao et al. (2024); Li et al. (2024). It is also challenging to perform a human-centered evaluation of the cultural nuances Liebling et al. (2022); Li et al. (2024). Existing studies have proposed strategies to evaluate cultural awareness of traditional MT systems by grounding images Khani et al. (2021), adapting entities Peskov et al. (2021) or targeting at dilates Riley et al. (2022). Different from existing culturally relevant MTs, we focus on evaluating the cultural awareness of MT by translating culture-specific items, a relatively underexplored area.
MT with Terminology
Previous studies on machine translation with terminology focused primarily on generic domains Dinu et al. (2019), or popular ones (e.g., law, medicine) Ghazvininejad et al. (2023). However, translating culture-specific items carries its own set of unique challenges because literal translations of CSIs may not exist in the target culture, making translation adaption crucial for target language readers to understand these terms Vinay and Darbelnet (1995). The adaptation can create semantic asymmetry between the source words and their translations, which makes traditional translation evaluation metrics focused on semantic alignment insufficient for cross-cultural translation Hershcovich et al. (2022).
External Knowledge for MT:
There have been multiple threads of research efforts on integrating external knowledge such as bilingual translation lexicons for neural machine translation systems, including probability interpolation of lexicons Arthur et al. (2016); Khandelwal et al. (2021), data augmentation by back-translations Hu et al. (2019), decoding with a phrase memory Wang et al. (2017), and pre-training with an entity-based denoising objective Hu et al. (2022). Despite their effectiveness, these methods require further fine-tuning of the original MTs. As the parameters of LLMs (e.g., ChatGPT) may not be accessible, we focus on tuning-free methods for integrating external knowledge in this study Dinu et al. (2019).
LLM-based MT:
Large language models, such as GPT-3 Brown et al. (2020), have proven effective in machine translation for various high-resource languages Hendy et al. (2023); Jiao et al. (2023). In particular, a few recent studies have investigated the performance of LLM-based MT, including formality control of translation outputs Garcia and Firat (2022), in-context translation ability during pre-training Shin et al. (2022), and multilingual translation Scao et al. (2022); Zhu et al. (2023). Moreover, previous work indicates that LLMs can integrate external knowledge in the context into translation Ghazvininejad et al. (2023); Li et al. (2024). However, the exploration of LLM-based MT on leveraging cultural knowledge to translate culture-specific items is still lacking.
3 CAMT Dataset
Figure 2: Overview of CAMT construction pipeline.
To minimize the need for human efforts while still obtaining diverse, high-quality CSI-centered translation pairs across multiple languages and cultures, we rely on Wikipedia to collect the data. The overall workflow of our data collection includes (1) building a wiki-centered cultural taxonomy (§3.1); (2) curating parallel sentences containing culturally-relevant entities (§3.2); and, (3) augmenting geo-metadata (§3.3). Figure 2 displays an overview of our data construction pipeline.
@@ 3.1 Cultural Taxonomy Extraction
Since culture is an abstract concept, it is difficult to capture fine-grained cultural characteristics from texts directly. With this consideration in mind, we referred to an existing CSI classification framework  Newmark (1988), which has been popularly used in the study of human translations of cultural concepts, to identify culturally relevant texts from Wikipedia. Specifically, there are five CSI categories in this framework, including: 1) ecology; 2) material culture; 3) social culture; 4) organizations, customs, ideas; 5) gestures and habits. Each entity-centered Wikipedia page is labeled by a variety of Wikipedia categories Asthana and Halfaker (2018). To save the efforts of matching each entity on Wikipedia with each CSI category, we map CSI categories with Wikipedia categories by manually creating a mapping table (in Table 9) to establish connections between the two categories. Ultimately, 18 Wikipedia categories are identified as culturally related. An entity is classified as culturally related if the category of its Wikipedia page maps to one of the CSI categories.
3.2 Culture Parallel Text Collection
To construct a culture parallel text corpus (e.g., for English-Chinese), we collect public text articles from Wikipedia’s translation tool that cover a wide range of cultural topics, and conduct sentence alignment to get parallel sentences (tools are detailed in Appendix §C). To expand the language coverage in our corpora, we also reuse open-source parallel corpora from OPUS Tiedemann (2016). These include Wikipedia v1.0 for English to French and Spanish, as well as Samanantar v0.2 for English to Hindi, Tamil, and Telugu. To identify sentences that contain culture-related items (Step $1$ in Figure 2), we first perform entity-linking  Ringgaard et al. (2017) to identify Wikipedia entities on the source texts, then use classification tool Asthana and Halfaker (2018) to classify the Wikipedia categories of these entities. The categories are further mapped to our CSI categories using the cultural taxonomy (§3.1). The mapping table is shown in Appendix B. Finally, we only keep sentences that contain entities belonging to CSI categories.
@@ 3.3 Cultural Knowledge Augmentation
Existing MT studies Arthur et al. (2016); Hu et al. (2022) have used external knowledge sources (e.g., Wikidata) to improve named entity translations. To enable future adaptations of these studies on our dataset, we parse Wikidata to augment cultural knowledge of CSIs (Step $2$ in Figure 2), which includes their translations, descriptions, and aliases in multiple languages. We also obtain the plain text of the first paragraph of the Wikipedia article as the CSI explanation. Moreover, to identify cultural items that are specific to a certain country, we collect information on country of origin from Wikidata for each item and remove sentences that contain only items without an associated country of origin. We refer to these groupings of data for each CSI as CSI dictionaries, an example of which is shown in the data example in Appendix §A. This meticulous approach enriches our dataset with supplementary cultural knowledge, enabling us to evaluate MTs’ performances in handling CSIs.
Dataset Language Domains Format
ApposCorpus en, es Person Sent.
Kementchedjhieva et al. (2020) de, pl Organization
Adaption en, de Celebrity Ent.
Peskov and Hangya (2021)
IdiomKB en, zh Idiom Sent.
Li et al. (2024) ja
CulturalRecipes en, zh Recipes Para.
Cao et al. (2024)
CAMT (Ours) en, zh, Cultural Sent.
fr, es Categories
hi, ta, te
Table 1: Dataset comparison. Sent., Ent., and Para. are abbreviations of sentence, entity and paragraph.
3.4 Dataset Analysis
We briefly compare CAMT with existing datasets that similarly focus on translating culture-specific content in Table 1. Similar to ApposCorpus and IdiomKB, translation pairs in CAMT are at the sentence level, aiming to provide fine-level textual context to explore the translation quality of CSIs from both semantic and pragmatic perspectives. Regarding data diversity, CAMT significantly expands the coverage to 18 cultural categories compared to prior work that focus on a specific domain. With respect to languages, CAMT includes 7 languages, which is more than in existing datasets ($\leq$ 4).
@@ 3.4 Dataset Analysis
We further conduct detailed corpus statistics on CAMT. As shown in Table 2, our dataset contains $6,948$ parallel sentences over $6$ language pairs, of which $3,029$ sentences have CSI translations and the rest are non-translation instances. The total number of unique CSIs (called CSIs Types) in CAMT is $6,983$. Among various cultural categories, we find that organizations, customs, ideas and material culture are the top 2 categories, and social culture and ecology are the bottom ones. A more detailed breakdown of statistics of CAMT can found in Appendix §D.
@@ 4 Cultural Awareness Evaluation
To better capture the cultural nuances in CSI translations, we devise two evaluation metrics: (1) CSI-Match, which evaluates the accuracy of CSIs with labels, and (2) PTA, which assesses the pragmatic translation quality of CSIs without labels.
@@
Existing evaluation metrics for terminology are efficient for evaluating the accuracy of translations and the fluency of outputs Anastasopoulos et al. (2021). However, previous metrics such as Exact Match (EM) assume that terminology translations must be exact matches, while reasonable adaptations of CSIs are also acceptable Vinay and Darbelnet (1995). To address this, we introduce the CSI-Match metric as a modification to the EM evaluation metric. CSI-Match measures the accuracy of term translation using a more nuanced, fuzzy matching approach. It calculates the maximal partial similarity ratio (PSR) between the reference CSI translations ${t_{1},t_{2},\ldots,t_{n}}$ and the system output sentence $S$. CSI-Match is determined by Eq. (1), resulting in a value from 0 to 100. A higher value indicates a stronger similarity of the predicted CSIs to a set of CSI translation references.
@@
where $S_{i:j}$ is a substring of $S$ from word index $i$ to $j$, and $d(,)$ is the normalized Levenshtein distance which measures the minimum number of insertions, deletions, and substitutions required to change one string into another  Levenshtein et al. (1966).
Pragmatic Translation Assessment (PTA):
Existing evaluation metrics like BLEU Papineni et al. (2002a) and COMET Rei et al. (2020a) mainly focus on surface-level or semantic-level accuracy between the source and target texts. However, in the context of CSI translation where translation quality is tightly associated with target culture Hershcovich et al. (2022), pragmatic translation quality becomes crucial. For example, a free translation based on the description of the CSI might have better pragmatic translation quality than a direct literal translation, as it would be easier for people from the target culture to understand. Therefore, we design a new assessment metric called PTA, measuring the win rate at which CSI translations by the MT system are judged to exhibit better pragmatic translation quality compared to human reference translations. Specifically, in the human evaluation, we specify what the CSIs are in the source language and ask native speakers of the target language to compare the CSI translations within the sentential context between an MT system and a reference, and then select the translation that is easier to understand. To improve the applicability of PTA when native speakers are not available, we use GPT-4o to replace human judgments for PTA, which has proven effective for the automatic evaluation of generative models in recent studies Rafailov et al. (2023); Kocmi and Federmann (2023). The evaluation prompt, shown in Appendix §E, is used for both human annotators and GPT-4. A higher PTA score means the system translates the CSI in a more comprehensive manner than the reference sentence does, which might use other accurate but less understandable translations of CSIs.
@@
cannoli:{CJK}UTF8gbsn里考塔芝士卷(Ricotta cheese rolls), 奶油甜馅煎饼卷 (Sweet Cream pancake rolls)
@@
Knowledge Translations:cannoli:{CJK}UTF8gbsn里考塔芝士卷(Ricotta cheese rolls), 奶油甜馅煎饼卷 (Sweet Cream pancake rolls)
@@
Table 3: Prompting strategy examples (Top) and a source with cultural knowledge for En-Zh translation (Bottom).